# PanoLAM: Large Avatar Model for Gaussian Full-Head Synthesis from One-shot Unposed Image

### v1

@@ Abstract
We present a feed-forward framework for Gaussian full-head synthesis from a single unposed image. Unlike previous work that relies on time-consuming GAN inversion and test-time optimization, our framework can reconstruct the Gaussian full-head model given a single unposed image in a single forward pass. This enables fast reconstruction and rendering during inference. To mitigate the lack of large-scale 3D head assets, we propose a large-scale synthetic dataset from trained 3D GANs and train our framework using only synthetic data. For efficient high-fidelity generation, we introduce a coarse-to-fine Gaussian head generation pipeline, where sparse points from the FLAME model interact with the image features by transformer blocks for feature extraction and coarse shape reconstruction, which are then densified for high-fidelity reconstruction. To fully leverage the prior knowledge residing in pretrained 3D GANs for effective reconstruction, we propose a dual-branch framework that effectively aggregates the structured spherical triplane feature and unstructured point-based features for more effective Gaussian head reconstruction. Experimental results show the effectiveness of our framework towards existing work.
@@ 1 Introduction
One line of work uses 3D Generative Adversarial Networks (3D-GANs) [11, 2] for the unconditional generation of 3D head avatars, which learn 3D generation models from 2D images paired with camera poses. Unlike 2D-GANs [33] for single-view generation, these approaches can generate face images with control over viewing angles. The pioneering work EG3D [11] introduces a triplane representation in Cartesian space to enable novel view synthesis after generation. The following works [2, 43] allow a wider range of viewing angles from the generated NeRF space and make 360 images synthesis possible. However, these GAN-based frameworks require time- and computation-consuming GAN-inversion [42] or test-time optimization [65] for image-conditioned generation during inference. Furthermore, the accurate camera pose of the input image is required for joint optimization of these inversion approaches. Rodin [77] and its follow-up [86] instead utilize diffusion models to generate triplanes of the head avatars. However, the multi-step diffusion process during inference still requires minutes of optimization for each case. Moreover, the costly ray marching hinders the rendering resolution of these NeRF-based frameworks. Though 2D super-resolution upsampling is proposed to improve the rendering efficiency and quality, it compromises 3D consistency.
@@
In this work, we propose a feed-forward framework to generate a full-head Gaussian avatar given a single unposed image. This enables fast reconstruction and rendering of a 3D avatar during inference. However, this task presents certain challenges. 1) The first hurdle lies in the lack of large-scale 3D head datasets for network training. To tackle this problem, we propose to leverage the prior knowledge in trained 3D GANs and sample a large-scale, diverse dataset from EG3D [11] and SphereHead [43]. We label and remove invalid cases manually to improve the quality of the dataset. 2) The second challenge involves how to efficiently and effectively reconstruct the high-fidelity Gaussian head from a single unposed image. To overcome this, we propose two main designs for our framework. i) Firstly, to improve the network efficiency, we propose a coarse-to-fine reconstruction mechanism for high-fidelity Gaussian full-head reconstruction. Specifically, sparse points extract rich image features, craft the coarse shape, and are then upsampled for high-fidelity dense Gaussian head reconstruction. By leveraging the topology prior in the FLAME model, we also introduce an efficient network-free upsampling mechanism for features and point clouds densification. ii) Secondly, to improve the effectiveness of the reconstruction process, we propose a dual-branch framework that hybridizes the unstructured point features and the structured spherical triplane features for the reconstruction of a Gaussian head. To distill the spherical triplane knowledge from the pretrained 3D GAN, we analyze its original feature aggregation manner and propose an efficient mechanism to effectively aggregate multi-layer features from the spherical triplane.
@@ 2 Related work
Various 3D representations have been applied for 3D head modeling, including mesh, Neural Radiance Fields (NeRFs) [54], Signed Distance Functions (SDFs), and Gaussian Splatting [35]. For mesh-based modeling, significant research [7, 61, 45] has been dedicated to representing 3D human heads with parametric textured meshes. However, predefined topologies constrain the capture of finer details and accessories such as hair and glasses, resulting in images that may appear unrealistic. More recent works [36, 48] try to improve the topology by deforming the FLAME model to cover these detailed areas. With the advent of NeRF for novel view synthesis, some recent works [83, 52, 46, 37, 4, 59, 95, 89, 90, 5, 27, 26, 60, 3, 31, 73, 25] also utilize NeRF for 3D head modeling and enable more realistic rendering. Along with NeRF, SDF, another implicit representation is also introduced into the 3D head modeling field [92, 91, 85, 9] to improve geometry quality, where each 3D position is assigned with the signed distance to the nearest surface. To enable more efficient rendering, the point-based [93] and Gaussian Splatting-based [63, 80, 66, 51, 23, 14] frameworks are also proposed. Although getting good performance, these 3D head modeling frameworks usually require minutes to hours of optimization from videos or multiview images with estimated camera pose for each person before usage, limiting their capability of scaling up and applications that require fast reconstruction. Instead, our work introduces a framework that can generate a Gaussian full-head avatar from a single unposed image in a single forward pass within seconds.
@@
One line of generative models utilizes 3D-aware Generative Adversarial Networks (GANs) [11] to synthesize view-consistent images by capturing the scene’s inherent 3D geometry and appearance from collections of training 2D images. Early approaches [56, 55, 24, 71, 68, 49] employ explicit 3D representations like meshes and voxel grids, while more recent studies [11, 2] utilize implicit representations for better image quality. EG3D [11] introduces a 3D triplane representation with GAN to generate heads that are capable of novel view synthesis. The following works [2, 43], further enable 360 novel-view synthesis. However, these GAN-based methods are mainly designed for unconditional generation, and require time- and computation-consuming techniques like GAN inversion [69, 12, 42] and test-time optimization [65] for image-conditioned generation. The slow optimization procedure during inference and the sacrifice in multi-view consistency hinder their real-world application.  [74, 6, 84, 13] distill an encoding network from EG3D for efficient image-to-triplane generation, but can only generate views near the front face. Instead of using GAN-based frameworks, more recent works [77, 86] utilize diffusion models to generate triplanes of head avatars. However, the multi-step diffusion process is still slow and computation-consuming during inference. Several works [17, 1, 70, 41, 22, 47] also introduce generalizable animatable avatar generation, but cannot reconstruct full heads. Moreover, due to the slow rendering speed of NeRF, these triplane NeRF-based approaches usually render low-resolution images from NeRF and apply super-resolution networks to enhance image detail, leading to view inconsistencies. To enhance rendering speed and rendered image quality, recent works [30, 16, 40, 72] also utilize Gaussian Splatting [35] as the 3D representation for generation. Specifically, LAM [30] introduces a feedforward framework for single-shot animatable Gaussian head generation that can be animated and rendered in real-time on various platforms. However, it requires estimation of FLAME parameters for reconstruction, which is time-consuming, and the generated head can only be rendered from limited viewing angles due to the limited viewing angles in the training 2D videos. This work, instead, can reconstruct the 3D Gaussian full-head given a single unposed image in a single forward pass efficiently, thanks to our large-scale 3D training datasets and our coarse-to-fine and dual-branch network design.
3D Avatar Datasets.
The quantity and quality of 3D training datasets are crucial to the capability and generalizability of learning-based models. Various datasets have been introduced into this field in the past decades. However, as shown in Table 1, many datasets [81, 88, 32, 67, 10, 15, 87, 19, 39, 18] only capture the front region of the faces. Networks trained on these datasets cannot model the side and back region of the head well. Although recent works [94, 76, 58, 29, 53] introduce 360 datasets into this field, the number of subjects in these datasets is limited, hindering the generalizability of trained networks. Under the observation that it’s much easier for humans to remove bad cases from generated 3D avatars compared to directly modeling them, we utilize pretrained 3D-aware GANs to develop and label a large-scale 3D head dataset. Leveraging two prior models, EG3D for 72 front views and SphereHead for 360 view synthesis, our total dataset, after removing bad cases manually to improve the quality, consists of 5M images of 165K subjects. To our knowledge, ours is the largest and most diverse dataset for 3D Avatar generation.

### v2

@@ Abstract
We present a feed-forward framework for Gaussian full-head synthesis from a single unposed image. Unlike previous work that relies on time-consuming GAN inversion and test-time optimization, our framework can reconstruct the Gaussian full-head model given a single unposed image in a single forward pass. This enables fast reconstruction and rendering during inference. To mitigate the lack of large-scale 3D head assets, we propose a large-scale synthetic dataset from trained 3D GANs and train our framework using only synthetic data. For efficient high-fidelity generation, we introduce a coarse-to-fine Gaussian head generation pipeline, where sparse points from the FLAME model interact with the image features by transformer blocks for feature extraction and coarse shape reconstruction, which are then densified for high-fidelity reconstruction. To fully leverage the prior knowledge residing in pretrained 3D GANs for effective reconstruction, we propose a dual-branch framework that effectively aggregates the structured spherical triplane feature and unstructured point-based features for more effective Gaussian head reconstruction. Experimental results show the effectiveness of our framework towards existing work. Project page at: https://panolam.github.io/.
Figure 1: PanoLAM creates high-fidelity Gaussian full-heads with one-shot unposed images in seconds.
@@ 1 Introduction
One line of work uses 3D Generative Adversarial Networks (3D-GANs) (Chan et al., 2022a; An et al., 2023) for the unconditional generation of 3D head avatars, which learn 3D generation models from 2D images paired with camera poses. Unlike 2D-GANs (Karras et al., 2019) for single-view generation, these approaches can generate face images with control over viewing angles. The pioneering work EG3D (Chan et al., 2022a) introduces a triplane representation in Cartesian space to enable novel view synthesis after generation. The following works (An et al., 2023; Li et al., 2024) allow a wider range of viewing angles from the generated NeRF space and make 360 images synthesis possible. However, these GAN-based frameworks require time- and computation-consuming GAN-inversion (Ko et al., 2023) or test-time optimization (Roich et al., 2023) for image-conditioned generation during inference. Furthermore, the accurate camera pose of the input image is required for joint optimization of these inversion approaches. Rodin (Wang et al., 2023a) and its follow-up (Zhang et al., 2024a) instead utilize diffusion models to generate triplanes of the head avatars. However, the multi-step diffusion process during inference still requires minutes of optimization for each case. Moreover, the costly ray marching hinders the rendering resolution of these NeRF-based frameworks. Though 2D super-resolution upsampling is proposed to improve the rendering efficiency and quality, it compromises 3D consistency.
@@
In this work, we propose a feed-forward framework to generate a full-head Gaussian avatar given a single unposed image. This enables fast reconstruction and rendering of a 3D avatar during inference. However, this task presents certain challenges. 1) The first hurdle lies in the lack of large-scale 3D head datasets for network training. To tackle this problem, we propose to leverage the prior knowledge in trained 3D GANs and sample a large-scale, diverse dataset from EG3D (Chan et al., 2022a) and SphereHead (Li et al., 2024). We label and remove invalid cases manually to improve the quality of the dataset. 2) The second challenge involves how to efficiently and effectively reconstruct the high-fidelity Gaussian head from a single unposed image. To overcome this, we propose two main designs for our framework. i) Firstly, to improve the network efficiency, we propose a coarse-to-fine reconstruction mechanism for high-fidelity Gaussian full-head reconstruction. Specifically, sparse points extract rich image features, craft the coarse shape, and are then upsampled for high-fidelity dense Gaussian head reconstruction. By leveraging the topology prior in the FLAME model, we also introduce an efficient network-free upsampling mechanism for features and point clouds densification. ii) Secondly, to improve the effectiveness of the reconstruction process, we propose a dual-branch framework that hybridizes the unstructured point features and the structured spherical triplane features for the reconstruction of a Gaussian head. To distill the spherical triplane knowledge from the pretrained 3D GAN, we analyze its original feature aggregation manner and propose an efficient mechanism to effectively aggregate multi-layer features from the spherical triplane.
@@ 2 Related Work
Various 3D representations have been applied for 3D head modeling, including mesh, Neural Radiance Fields (NeRFs) (Mildenhall et al., 2020), Signed Distance Functions (SDFs), and Gaussian Splatting (Kerbl et al., 2023). Mesh-based modelings (Blanz & Vetter, 1999; Paysan et al., 2009; Li et al., 2017) are dedicated to representing 3D human heads with parametric textured meshes and deform the mesh model to improve the topology (Khakhulin et al., 2022; Liao et al., 2025). NeRF-based works (Yu et al., 2023; Ma et al., 2023; Li et al., 2023a; Ki et al., 2024; Bai et al., 2023a; Park et al., 2021a; Zielonka et al., 2023; Zhang et al., 2024b; Zhao et al., 2024; Bai et al., 2023b; Guo et al., 2021; Gao et al., 2022; Park et al., 2021b; Athar et al., 2022; Hong et al., 2021; Tretschk et al., 2021; Gafni et al., 2021) utilize NeRF for 3D head modeling and get more realistic rendering. To improve geometry quality, SDF-based methods (Zheng et al., 2025; 2022; Zanfir et al., 2024; Canela et al., 2024) are also introduced, where each 3D position is assigned with the signed distance to the nearest surface. For more efficient rendering, the point-based (Zheng et al., 2023) and Gaussian Splatting-based (Qian et al., 2024; Xu et al., 2024; Saito et al., 2024; Ma et al., 2024; Dhamo et al., 2024; Chen et al., 2024) frameworks are also proposed. However, these frameworks usually require minutes to hours of optimization from videos or multiview images with estimated camera pose for each person before usage, limiting their capability of scaling up and applications that require fast reconstruction. Instead, our work introduces a framework that can generate a Gaussian full-head avatar from a single unposed image in a single forward pass within seconds.
@@
One line of generative models utilizes 3D-aware Generative Adversarial Networks (GANs) (Chan et al., 2022a) to synthesize view-consistent images. Early approaches (Nguyen-Phuoc et al., 2020; 2019; Gadelha et al., 2017; Szabó et al., 2019; Shi et al., 2021; Liao et al., 2020) employ explicit 3D representations like meshes and voxel grids, while more recent studies (Chan et al., 2022a; An et al., 2023) utilize implicit representations for better image quality. EG3D (Chan et al., 2022a) and its follow-ups (An et al., 2023; Li et al., 2024) utilize 3D triplane representation with GANs to generate heads that are capable of novel view synthesis. However, these methods are mainly designed for unconditional generation, and require time- and computation-consuming techniques like GAN inversion (Sun et al., 2022; Chan et al., 2022b; Ko et al., 2023) and test-time optimization (Roich et al., 2023) for image-conditioned generation. The slow optimization procedure during inference and the sacrifice in multi-view consistency hinder their real-world application.  Trevithick et al. (2023); Bhattarai et al. (2024); Yuan et al. (2023); Chen et al. (2023) distill an encoding network from EG3D for efficient image-to-triplane generation, but can only generate views near the front face. Instead of using GAN-based frameworks, more recent works (Wang et al., 2023a; Zhang et al., 2024a) utilize diffusion models to generate triplanes of head avatars. However, the multi-step diffusion process is still slow and computation-intensive during inference. Several works (Chu et al., 2024; Abdal et al., 2024; Sun et al., 2023; Kirschstein et al., 2025; Deng et al., 2024; Li et al., 2023b) also introduce generalizable animatable avatar generation, but cannot reconstruct full heads. Moreover, due to the slow rendering speed of NeRF, these approaches require 2D super-resolution networks to enhance image detail, causing view inconsistencies. To enhance rendering quality and speed, recent works (He et al., 2025; Chu & Harada, 2024; Kirschstein et al., 2024; Tang et al., 2024) also utilize Gaussian Splatting (Kerbl et al., 2023) as the 3D representation for generation. LAM (He et al., 2025) introduces a feedforward framework for single-shot animatable Gaussian head generation that can be animated and rendered in real-time on various platforms. However, it requires estimation of FLAME parameters for reconstruction, which is time-consuming, and the generated head can only be rendered from limited viewing angles due to the limited viewing angles in the training 2D videos. Instead, this work can reconstruct the 3D Gaussian full-head given a single unposed image in a single forward pass, thanks to our large-scale 3D training datasets and our coarse-to-fine and dual-branch network design.
Figure 2: Overall Framework. Given an unposed head image as input, PanoLAM involves two branches to achieve single-pass 3D Gaussian head reconstruction: a point-based transformer for coarse-to-fine point shape reconstruction and point features extraction, and a spherical triplane transformer to distill prior knowledge from 3D GAN. Features from the two branches are concatenated for high-fidelity Gaussian head regression.
3 Methodology
3.1 Overview
As shown in Fig. 2, given an unposed reference image, our framework generates the Gaussian full-head in an efficient coarse-to-fine and joint point-triplane manner, where dense point cloud and point features are densified from sparse ones and fused with the spherical triplane features for dense Gaussian head regression. As shown in the upper left part of Fig. 2, in the coarse stage of the point branch, sparse vertices from the FLAME vertices interact with the extracted multi-level image features by stacked cross-attention modules for feature extraction and coarse Gaussian point reconstruction. The dense reconstruction stage leverages barycentric weights from FLAME subdivision to densify the coarse Gaussian point position and features for dense Gaussian points regression. To fully leverage the prior knowledge residing in the spherical triplane of 3D GAN, as shown in the bottom branch of Fig. 2, a spherical triplane branch is also introduced, and a novel efficient multi-layer aggregation mechanism is proposed to aggregate the structured features from the spherical triplane using the densified point cloud, which are then concatenated with the unstructured point features for high-fidelity Gaussian head reconstruction. To optimize the spherical-triplane branch and distill the prior knowledge, in the bottom right of the framework, a frozen SphereHead decoder is utilized to decode the feature to images for supervision during training.
3.2 Coarse-to-fine Gaussian Head Avatar Generation
The number of Gaussian points is one key factor that can affect the fidelity of the reconstructed Gaussian head, especially for the head avatar that contains sharp details like hair strands, mustaches, and wrinkles. The total number of Gaussian points should be large enough to produce these details. However, the large number of Gaussian points is hard to optimize during training, especially in our feed-forward pipeline, where the iterative densification operation in the traditional 3D Gaussian Splatting training pipeline is not applicable. To resolve this problem, unlike previous works (He et al., 2025; Zou et al., 2024) that directly regresses the dense point cloud, we introduce a coarse-to-fine mechanism. Specifically, sparse Gaussian points are first reconstructed for coarse shape reconstruction and then densified to dense Gaussians for high-fidelity details crafting. There are two advantages in the proposed design. Firstly, the sparse number of points enables efficient cross-attention with the image features for textures and shape features extraction, saving large computation and memory costs compared with a large number of points. Secondly, the sparse point can be well supervised and optimized for large deformation to carve the shape of the head, especially in regions that require large deformation from the original FLAME model, e.g., long hair. Dense Gaussian points upsampled from these correctly deformed sparse points can focus more on texture details crafting, which are much easier to optimize.
Coarse Gaussian Head Generation
To fully leverage the head shape prior that resides in the FLAME model, we initialize the coarse Gaussian point with sparse vertices (5,023 points) on the neutral FLAME model in the canonical space. Our target is to regress the deformation offset for shape refinement of regions that FLAME cannot model, e.g., long hairs, glasses, and caps. To supervise the shape deformation with RGB input images, Gaussian attributes for each point are also regressed for differentiable Gaussian Splatting. To get started, we assign positional embedding to each point and utilize learnable multi-layer perceptron (MLP) modules to project the channel of features into the token channel $C_{t}$ of transformers as: $F_{P_{0}}=MLP(\gamma(V)),$ where $V$ is the spatial position of each vertice and $\gamma$ the L_\_frequency sinusoidal encoding as in NeRF. To extract the texture and shape information in the given unposed image, we utilize the pre-trained DinoV2 (Oquab et al., 2024) for feature extraction. Inspired by  (Ranftl et al., 2021; He et al., 2025), we fuse features derived from both shallow and deep layers to obtain both local and global image features $F_{I}$. Specifically, an MLP fuses the $\{5,12,18,24\}$ layers of features in DinoV2 into $C_{t}$ token channels of features as:
$F_{I}=MLP(\mathcal{C}(F_{D_{5}},F_{D_{12}},F_{D_{18}},F_{D_{24}}),$ (1)
Given mapped point features $F_{P}$ and extracted image features $F_{I}$, we utilize $L_{\mathcal{A}}$ layers of stacked cross-attention modules $\mathcal{A}=\{\mathcal{A}_{i}\}_{i=1}^{L_{\mathcal{A}}}$ from Transformer (Vaswani et al., 2017) for feature extraction from image to point cloud as follows:
$F_{P_{i}}=\mathcal{A}_{i}(F_{P_{i-1}},F_{I}),$ (2)
$\{c_{k},o_{k},s_{k},R_{k},O_{k}\}^{M_{C}}_{k=1}=\mathcal{D}(F_{P_{k}}),$ (3)
Efficient Point Cloud Densification
After the reconstruction of the coarse Gaussian head, we obtain a sparse point cloud that describes the shape of the target head avatar in the image, each point is also attached with rich features extracted from the image. We then densify these points and features for high-fidelity dense Gaussian head reconstruction. Unlike previous methods (Li et al., 2021; Yu et al., 2018) that require a network trained to upsample sparse points, we propose to leverage the topology in the FLAME mesh for efficient densification. Specifically, the proposed strategy utilizes the barycentric coordinates from mesh subdivision to densify the point cloud and feature. Our key observation is that after the deformation in the coarse reconstruction stage, the topology between each point mostly remains as the original FLAME vertices. Thereafter, we can precompute the barycentric coordinates for point cloud densification utilizing mesh subdivision on FLAME. Specifically, given the original FLAME model, we perform subdivision for faces that are larger than a threshold $\lambda_{area}$, the position of newly added vertices $v^{\prime}$ can be computed from the three vertices of the original face with barycentric coordinates as: $v^{\prime}=\sum_{i=1}^{3}({w_{i}\cdot v_{i}}),$ where $w_{i}$ and $v_{i}$ are the barycentric weight and vertex coordinate of the original face, respectively. After the deformation in the coarse reconstruction stage, each vertex moves to a new position w.r.t the predicted offset as $p_{i}=v_{i}+O_{i}$. Then new added point $p^{\prime}$ with its attached point features $p_{F}^{\prime}$ can be interpolated with barycentric weights as: $p^{\prime}=\sum_{i=1}^{3}({w_{i}\cdot p_{i}})$, and $p_{F}^{\prime}=\sum_{i=1}^{3}({w_{i}\cdot p_{F_{i}}}),$ where $p_{F_{i}}$ is the extracted features from the last cross-attention blocks in Equ. (2). In this way, we can efficiently densify the point and features.
Dense Gaussian Head Generation.
After getting the dense point position with dense point features, we can utilize them for dense Gaussian head regression. Rather than using only unstructured point features for prediction, we also aggregate multi-layer of features from the structured spherical triplane for more effective Gaussian attributes regression, which we will discuss in Sec. 3.3. Denote $F_{P_{i}}$ the $i_{th}$ point feature and $F_{T_{i}}$ the aggregated spherical triplane feature for point $P_{i}$, the dense Gaussian head attributes and a small deformation residual can be regressed similarly to the coarse one as:
3.3 Spherical Triplane Prior Knowledge Distillation
In the previous section, we obtain the dense point cloud with its attached point features extracted from the unposed image. To better leverage the prior knowledge within SphereHead’s framework, we introduce a spherical triplane branch to distill the knowledge from a pretrained model.
Preliminary: The Spherical Triplane representation
from SphereHead Li et al. (2024) combines the three triplanes in the Cartesian coordinate system with another three spherical planes in the spherical coordinate system, where features are queried with projection operations in the respective coordinate systems. This representation alleviates the feature entanglement issue in regular triplanes.
In the spherical triplane branch, we initialize the $H_{T}\times W_{T}\times 6\times C_{t}$ spherical triplane features with learnable query features, which are then flattened into token features $F_{T}\in\mathbb{R}^{N_{T}\times C_{t}}$, where $N_{T}=H_{T}\times W_{T}\times 6$. These learnable features then interact with the image features $F_{I}$ with stacked cross-attention blocks $\mathcal{A_{T}}=\{\mathcal{A_{T}}_{i}\}_{i=1}^{L_{\mathcal{A_{T}}}}$ similar to the learnable point features as:
$F_{T_{i}}=\mathcal{A_{T}}_{i}(F_{T_{i-1}},F_{I}),$ (5)
where $F_{T_{i}}$ is the $i_{th}$ layer of spherical triplane features in the stacked cross-attention layers. The final output features are then projected by MLPs to the same channel as the spherical triplane in SphereHead. We then need to query the feature from the spherical triplane to enhance the point features. One vanilla way is to utilize the point-based query strategy that fetches each point’s spherical-triplane feature by projection. However, we find that our point clouds are distributed on the shape surface similar to a mesh, and such a single-layer query strategy cannot fully aggregate the corresponding features of each point from the spherical triplane.
That’s because the triplane designed for NeRF-based rendering requires ray marching to aggregate multiple points(/layers) of features for rendering. Using a single point cloud on the shape surface cannot fully fetch valuable features from the spherical triplane. While ray marching is computationally consuming and inefficient, we propose to sample multiple layers of points from the single-layer point cloud for feature aggregation. As is shown in Fig. 3, we sample 4 virtual cameras that can capture the front, left, right, and back sides of the head. For each camera, we cast a ray from the camera center to each point $P_{i}$, then $K_{m}$ number of points $P^{\prime}=\{P^{\prime}_{i}\}_{i=1}^{K_{m}}$ are sampled along the ray near the point. Each point then queries features from the spherical triplane by projection. We then concatenate and aggregate these queried features with MLPs to obtain the final spherical triplane features as:
$F_{T_{i}}=MLP(\mathcal{C}(\{F^{\prime}_{T_{i}}\}_{i=1}^{K_{m}})),$ (6)
where $F^{\prime}_{T_{i}}$ is the spherical triplane feature queried by sampled point $P^{\prime}_{i}$ ; $\mathcal{C}$ is the concatenation operation; and $F_{T_{i}}$ is the final aggregated feature for point $P_{i}$. In this way, each point gets its aggregated spherical triplane feature. These aggregated spherical triplane features can be concatenated with the point feature for dense Gaussian attribute regression as in Formula (4). To supervise this branch in distilling knowledge from SphereHead, we add a frozen decoder from SphereHead to decode the feature into the target image, supervised by ground truth images during training, as shown in the bottom right part of Fig 2.
3.4 Large-scale Training Data Generation